HUBBLE SPACE TELESCOPE DAILY REPORT #1649 PERIOD COVERED: 0000Z (UTC) 06/05/96 - 0000Z (UTC) 06/06/96 Daily Status Report as of 158/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOC/96 6349 (Snapshot search for High Redshift QSOs with = Far-UV Flux) The Faint Object Camera (f/96) was used to make snapshot images of = the quasi-stellar objects Q1251+367 and Q0201-504. This is to determine = if these high redshifted objects have a far ultraviolet flux. The = observations were executed as scheduled, and no anomalies were noted. 1.2 Completed WF/PC-2 6253 (Pure Parallel Archive Program) The WF/PC-2 was used to execute a pure parallel program taking = ultraviolet observations running concurrently with prime science = observations. The intent is to search for extra-galactic sources with = flux excesses at 3000 A0 and galactic sources with excesses at 1600 A0. = The observations were completed as planned, and no problems were = reported. 1.3 Completed WF/PC-2 5997 (The Cores of Elliptical Galaxies in the = Coma Cluster) The WF/PC-2 was used to make observations of the large elliptical = galaxies IC3947 and Gal-125930 which are located in the Coma Cluster. = The purpose is to study the luminosity profiles in their central nuclear = region. The observations were completed as scheduled, and no anomalies = were noted. 1.4 Completed Two Sets of WF/PC-2 6361 (High Resolution Imaging of = X-Ray Selected AGNs) The WF/PC-2 was used to make high resolution observations of the = known low red shifted x-ray active galactic nuclus (AGN) MS0721.2+6. = These data will allow a study of the optical properties of an unbiased = sample of AGN host galaxies. The observations were completed as = scheduled, and no problems were reported. 1.5 Completed GHRS/2/1 5830 (The Small Icy Particles of the A and B = Rinds of Saturn) The High Resolution Spectrograph (Sides One and Two) were used to = make spectrographic observations of the A and B rinds on the planet = Saturn. Voyager and IUE spectra show a strong absorption near 1650 A0 = that may be to small water ice particles and/or water vapor. The = observations were executed as planned and no anomalies were reported. 1.6 Completed WF/PC-2 6189 (WF/PC-2 Cycle 5 Calibration: Internal = Monitor - Vis Flats) The WF/PC-2 was used to make internal visual flat calibration = images. Both photometric and narrow-band filters were used. The = observations were executed as planned, and no problems were reported. 1.7 Completed WF/PC-2 6116 (A Search For the Optical Counterpart of a = Triple Pulsar System PRS-B1620) The WF/PC-2 was used to make observations to search for the = optical counterpart of the triple pulsar system PSR-B1620. The = observations were executed as scheduled, and no anomalies were noted. 1.8 Completed FOS/RD 6577 (An Expanded Survey for Study of the = Lyman-Alpha Line and QSO Mg II Absorption) The Faint Object Spectrograph (Red Side) was used to make = observations of the QSO UM366 in order to make a survey of the = Lyman-Alpha line and the 68 Mg II absorption line system. The = observations were completed as scheduled, and no problems were reported. 1.9 Completed WF/PC-2 5917 (The Globular Clusters in the Fornax Dwarf = - Physical Parameters) The WF/PC-2 was used to search for globular clusters in dwarf = galaxy Fornax-023 which is a member of the Fornax Galactic Cluster. The = cluster age, metal abundances, and radial profiles will be determined = from the data. Because the galaxies in the Fornax Cluster have different = morphologies than those of the Local Group, the globular cluster systems = are also expected to show differences. The observations were executed as = scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 Scheduled Reacquisitions : 10 Successful : 10 2.2 FHST Updates: Scheduled : 15 Successful : 15 2.3 Operations Notes: HST operations reports no anomalous activity during the last 24 = hours. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH